# Graph sub-sampling for divide-and-conquer algorithms in large networks

### v2

@@ 2.2.2 Divide-and-conquer algorithm
With a metric to quantify the CP structure of a network, we can present the divide-and-conquer algorithm. We slightly modify the algorithm in Yanchenko, (2022) for reasons that will be plain later. The idea is to find the CP structure on small sub-networks of the network and then combine these results together to yield the CP labels on the entire network. Specifically, let $A$ be the adjacency matrix with $n$ nodes, $q\in(0,1)$ be the proportion of nodes to sub-sample and $B$ the number of sub-sample to draw. First, we sample a sub-graph with $qn$ nodes using some graph sub-sampling routine, and find the optimal CP labels from (3). We repeat this process $B$ times and output $\hat{\boldsymbol{c}}\in[0,1]^{n}$ where $\hat{c}_{i}$ is the number of times node $i$ was assigned to the core over the total number of samples $B$. The full algorithm is presented in Algorithm 2.
@@
Input: $n\times n$ adjacency matrix $A$, proportion of nodes to sub-sample $q$, number of sub-samples drawn $B$, sub-sampling scheme $\mathcal{S}$
Initialize ${\boldsymbol{x}}={\bf 0}_{n}$ where $x_{i}$ is the number of times node $i$ is assigned to the core for $i=1,\dots,n$
@@
       Randomly sample sub-network $S^{(b)}$ using $\mathcal{S}$ with nodes $V=\{v_{1},\dots,v_{qn}\}$  Obtain CP labels of $S^{(b)}$, $\hat{\boldsymbol{c}}^{(b)}\in\{0,1\}^{qn}$$x_{v_{i}}=x_{v_{i}}+1$ if $\hat{c}^{(b)}_{i}=1$
@@
$\hat{c}_{i}=x_{i}/B=$ of times node $i$ was assigned to core / # of sub-samples
@@
This algorithm is slightly different from the original paper. Originally, the output was the number of times node $i$ was assigned to the core, divided by the number of times that node $i$ was sampled. Indeed, this approach more closely resembles the stitching routine of the $\mathsf{PACE}$ algorithm. The reason that $\mathsf{PACE}$ divides by the number of times that nodes $i$ and $j$ were sub-sampled together, $N_{ij}$, and then only keeps the estimate if $N_{ij}>\beta$ is that the estimate $\hat{C}_{ij}$ would be unreliable if it was based only on a few sub-samples. For the CP task, however, if a node is sub-sampled infrequently, this gives us more information, rather than less. Indeed, core nodes are those that are more central or influential in the network. Therefore, if a node is sampled infrequently, then it is unlikely to be important and also unlikely to be a part of the core. Thus, dividing by the total number of samples will decrease the estimate of $\hat{c}_{i}$ for infrequently sampled nodes. Implicitly, this means we should favor sub-sampling methods which have a high probability of sampling core nodes. As we will see, this small change not only makes proving theoretical statements easier, but also improves the empirical performance of the algorithm.
@@ 2.3 Sub-graph sampling
The simplest sampling method is to randomly sample nodes without replacement. Each node has an equal probability of being selected, and once all $qn$ nodes are chosen, then the edges between these nodes make up the sub-graph. This method was originally considered in both Mukherjee et al., (2021) and Yanchenko, (2022).
@@
In this method, nodes are again sampled at random, but now their probability of being sampled is proportional to their degree. Once the $qn$ nodes are sampled, then the sub-graph is completed with all associated edges between sampled nodes. Mathematically, this means sampling without replacement from a distribution with probability mass function $\mathsf{P}(Z=j)=d_{j}/\sum_{k}d_{k}$ where $d_{j}$ is the degree of node $j$.
@@
Instead of randomly sampling nodes, this approach samples edges uniformly at random. Edges are drawn with equal probability and the two nodes connected by the sampled edge are added to the sub-graph. This process continues until $qn$ nodes have been sampled, at which point the remaining edges between sampled nodes are also added. Adding the remaining edges between sampled nodes is sometimes called the Induction Step (Ahmed et al.,, 2013) and will be used in all remaining algorithms as well. While edges can be selected with equal probability, if the graph has edge weights than this could be used to for a sampling scheme analogous to DN.
@@
The previous three methods randomly sampled nodes or edges. For the remaining methods, we will consider graph exploration approaches. The BFS algorithm begins by sampling a node at random. Then all neighbors of this sampled node are included in the sub-graph as well. Next, all neighbors of the neighbors are also included in the sub-sample. This process continues until we have sampled $qn$ nodes.
@@
DFS can be considered as the complement of a BFS algorithm. This method also begins by randomly sampling a starting node. Then instead of including all neighbors of this starting node, it randomly chooses one neighbor. From there, it continues down this “branch” of the network. The process continues until $qn$ nodes are sampled, or the search terminates as there are no more nodes to traverse. If the latter happens, a new node is sampled to start the process again until $qn$ nodes have been reached.
@@
This approach shares similarities to BFS. As in BFS, a node is sampled at random and then all of its neighbors as also added to the sub-graph. Instead of looking at the neighbors of neighbors as in BFS, RNN then samples a new node and includes all of the new node’s neighbors. This continues until $qn$ nodes have been sampled.
@@
In our final algorithm, we traverse the network through a random walk. Specifically, we begin by sampling a node uniformly at random. Then from this node, one of its neighbors is sampled at random, and this process continues until the walk has traversed $qn$ nodes.
@@ 3 Theoretical Results
In this section, we present theoretical results for each algorithm. As this paper’s focus is on the sub-sampling step, we will primarily focus on the role of the sampling scheme on the theoretical results. Additionally, Mukherjee et al., (2021) devoted significant attention to the theoretical properties of $\mathsf{PACE}$, so we only briefly discuss these before devoting the majority of the space to new theoretical results on the CP divide-and-conquer algorithm.
@@ 3.1 $\mathsf{PACE}$ algorithm
The goal of the theoretical results for $\mathsf{PACE}$ is to show that the estimated clustering matrix $\hat{C}$, is “close” to the true clustering matrix $C$ which generated the network. To formalize this notion, let $\tilde{\delta}(C^{*},\hat{C})$ be the difference between the true matrix $C^{*}$ and estimate $\hat{C}$ where
$\tilde{\delta}(C,C^{\prime})=\frac{1}{n^{2}}||C-C^{\prime}||_{F}^{2}$
and $||\cdot||_{F}$ is the Frobenius norm. Thus, $\tilde{\delta}(C,\hat{C})$ is the mis-clustering rate of the $\mathsf{PACE}$ algorithm. In the following theorem, the authors bound the expectation of this term under a general community detection algorithm and sampling scheme.
Theorem 3.1 from Mukherjee et al., (2021): Given some network $A$ with $K$ communities where the largest community has $p_{max}n$ nodes, and some sampling scheme $\mathcal{S}$, let $S$ be a randomly chosen sub-graph of the network. Moroever, let $\hat{C}$ be the estimated clustering matrix returned by $\mathsf{PACE}$. Then the expected mis-clustering rate, $\mathbb{E}\tilde{\delta}(C,\hat{C})$ can be bounded by:
$\mathbb{E}\tilde{\delta}(C^{*},\hat{C})\leq\frac{B}{\beta n^{2}}\mathbb{E}||\hat{C}^{(S)}-C^{(S)}||^{2}_{F}+p_{max}\max_{i,j}\mathbb{P}(N_{ij}<\beta)$ (4)
where $\hat{C}^{(S)}$ and $(C^{*})^{(S)}$ are the estimated and true clustering matrices, respectively, restricted to sub-graph $S$.
See Mukherjee et al., (2021) for the proof of this result. There are two main parts to the bound in (4). The first term quantifies the performance of the community detection algorithm on a randomly chosen sub-graph $S$ which is mainly a function of the clustering algorithm. On the other hand, the second term measures how well the graph is covered by the sub-samples, and depends only on the sub-sampling procedure.
As our interest is on the sub-sampling step, we focus on this second term. Indeed, the authors of $\mathsf{PACE}$ provide two special cases of the results under the RN sampling scheme and Ego neighborhoods. As we do not discuss ego neighborhoods, we omit discussion of this result. The following gives the results for RN sampling.
Corollary 3.1 from Mukherjee et al., (2021): Assume that nodes are chosen uniformly at random to construct the subgraphs. Then
@@
This result shows that the mis-clustering error coming from the sub-sampling step decreases exponentially with $B$, the total number of sub-samples drawn, as well as $q$, the proportion of nodes in each sub-graph. As the authors note, deriving the specific result for other sub-sampling routines is difficult, so we do not consider them here.
@@ 3.2 CP algorithm
Using similar ideas to that of Mukherjee et al., (2021), we present novel theoretical results for the CP divide-and-conquer algorithm. Similar to the previous sub-section, we are interested in bounding the expected mis-classification error.
@@ 3.2.1 General mis-classification result
First, we derive the expression for the general mis-classification rate. Recall from Algorithm 2 our core estimates $\hat{\boldsymbol{c}}$, where
@@
and $\hat{\boldsymbol{c}}^{(b)}$ are the CP labels from sub-sample $b$, i.e., $\hat{c}_{i}^{(b)}=1$ if node $i$ was assigned to the core in sub-sample $b$, and 0 otherwise. Then we define
@@
to quantify the mis-clustering of the algorithm. The following result is analogous to Theorem 3.1 from Mukherjee et al., (2021).
Theorem 3.2: Given some network $A$ and sampling scheme $\mathcal{S}$, let $S$ be a randomly chosen sub-graph. Let $\hat{\boldsymbol{c}}$ be estimated CP labels using the CP divide-and-conquer algorithm. Then the expected mis-clustering rate, $\mathbb{E}\delta(\hat{\boldsymbol{c}},\boldsymbol{c}^{*})$ can be bounded by
@@
Please see the Supplemental Materials for all proofs in this sub-section. Similar to the results of $\mathsf{PACE}$, our error bound has two terms; the first which primarily depends on the CP algorithm, and the second which only depends on the sub-sampling routine. Indeed, if the expected fraction of mis-classified nodes and the expected fraction of core nodes that are not sampled both go to 0 for large $n$, then this entire term will converge to 0.
@@ 3.2.2 Specific sub-sampling results
The results of Theorem 3.2 are agnostic to the network-generating model, CP detection algorithm, and sub-sampling scheme. By assuming a particular model, we can find the analytic expression for the second term in (6) for the majority of the sub-sampling schemes in Section 2.3.
@@
We assume networks are generated with a CP structure from the stochastic block model (SBM) (Holland et al.,, 1983). For adjacency matrix $A$, let $A_{ij}\stackrel{{\scriptstyle ind.}}{{\sim}}\mathsf{Bernoulli}(\varrho_{n}P_{ij})$ where $P$ corresponds to a SBM that does not depend on $n$, and $\varrho_{n}\to 0$ is a sparsity-inducing parameter. Furthermore, if $\boldsymbol{c}^{*}$ correspond to the true CP labels, then
$P_{ij}=\begin{cases}p_{11}&c_{i}^{*}=c_{j}^{*}=1\\ p_{22}&c_{i}^{*}=c_{j}^{*}=0\\ p_{12}&\text{otherwise}\\ \end{cases}.$
@@
If $p_{11}>p_{12}>p_{22}$, then this model generates networks with a CP structure, so we call it the CP-SBM. We may also assume that $\alpha_{n}=k/n$, the proportion of core nodes, goes to zero to model the empirical observation that the core is typically smaller than the periphery.
In the following Corollary, we find an analytic expression for the expected number of core nodes that are not sampled in each sub-graph. We assume that $n$ is large enough such that the probability of sampling the same node twice is negligible. The result is in Corollary 3.2 with proof in the Supplemental Materials.
@@
Corollary 3.2: Let $A$ be generated from a CP-SBM with $n$ nodes and $k$ core nodes. Let $S$ be a sub-graph of $A$ of size $qn$ where $0\leq q\leq 1$ sampled using $\mathcal{S}$ where $\mathcal{S}$ is one of the sub-sampling schemes from this paper. If $y_{i}^{(S)}$ is the event that node $i$ is a core node and not sampled in $S$, then
$\mathbb{E}\frac{1}{n}\sum_{i=1}^{n}y_{i}^{(S)}=(1-q\xi_{n}^{(\mathcal{S})})\frac{k}{n}$ (7)
where $\xi_{n}^{(\mathcal{S})}$ takes the expression listed in Table 1.
$\mathcal{S}$ $\xi_{n}^{(\mathcal{S})}$ $\lim_{n\to\infty}\xi_{n}^{(\mathcal{S})}$
RN 1 1
DN $\frac{n\{(k-1)p_{11}+(n-k)p_{12}\}}{k(k-1)p_{11}+2k(n-k)p_{12}+(n-k)(n-k-1)p_{22}}$ $\frac{p_{12}}{p_{22}}$
RE $\frac{n\{(k-1)p_{11}+(n-k)p_{12}\}}{k(k-1)p_{11}+2k(n-k)p_{12}+(n-k)(n-k-1)p_{22}}$ $\frac{p_{12}}{p_{22}}$
RNN $\frac{1+(k-1)p_{11}+(n-k)p_{12}}{1+\frac{k}{n}\{(k-1)p_{11}+(n-k)p_{12}\}+\frac{n-k}{n}\{kp_{12}+(n-k-1)p_{22}\}}$ $\frac{p_{12}}{\alpha_{n}p_{12}+p_{22}}$
RW * *
Table 1: Theoretical results for sub-sampling scheme in CP divide-and-conquer algorithm.
The expression for RW (*) appear in the Supplemental Materials.
The proof carefully computes the expected number of cores nodes that are sampled in each sub-graph. For RN, DN and RE, we express the number of core nodes sampled as a hyper-geometric distribution. For RNN, the number of core nodes sampled is conditional on whether the first node drawn is from the core or periphery, so we consider these two cases. Lastly, we use a recursive relationship to model the RW case.
@@
In (7), we see the general result for the average number of core nodes that are not sampled in a given sub-graph. As expected, as size of the sub-graphs increases ($q$), this term decreases. Moreover, this expression depends on $k/n=\alpha_{n}$. Assuming that the periphery dominates the core, this entire term will go to zero regardless of the sampling scheme which theoretically validates our decision to divide the core estimates by $B$. Of course, the sub-sampling routine will still affect the performance for finite $n$.
@@
In Table 1 we present the form of the expression in (7) for five of the seven routines from Section 2.3. From (7), the overall error is smaller when $\xi_{n}^{(\mathcal{S})}$ is larger so large values of $\xi_{n}^{(S)}$ are to be preferred. BFS and DFS are difficult to analyze so these are omitted, and RW has a complicated expression that we relegate to the Supplemental Materials. For the other sub-sampling schemes, we first notice that DN and RE yield identical expressions. This means that these sampling schemes are equivalent in terms of the number of core nodes that are expected to sample in each sub-graph when the network is generated from a CP-SBM. This does not mean that they will perform identically as the sub-sampling scheme affects the topology of the sub-graph $S$ which may affect the first term in (6). Next, the exact expressions for $\xi_{n}^{(\mathcal{S})}$ are a bit cumbersome, so it is helpful to look at the asymptotic expression. Indeed, for DN and RE, this becomes $p_{12}/p_{22}$, the ratio of the core-periphery edge probability to the periphery-periphery edge probability. In a CP-SBM, we have $p_{12}>p_{22}$ so $p_{12}/p_{22}>1$. This means that DN and RE have a uniformly smaller expected error than RN. Additionally, as the strength of the CP structure of the network increases, the error will decrease because $p_{12}/p_{22}$ increases. In other words, we expect a lower error rate if the network exhibits a stronger CP structure. Finally, for the asymptotic results of RNN, $p_{12}/(\alpha_{n}p_{12}+p_{22}<p_{12}<p_{22}$. Thus, DN and RE are expected to have a lower error than that of RNN. Moreover, if $(1-\alpha_{n})p_{12}<p_{22}$, then RNN is also expected to have a worse error rate than that of RN.
@@
It is important to highlight that these results and discussion apply only to the second term in (6). It is possible that a sub-sampling method may not sample core nodes as often, but has a lower mis-classification rate on the sub-graph, leading to a smaller value of the first term in (6), or vice-versa. Studying this term is more difficult, especially for methods besides RN, so we leave this as an avenue of future work.

### v3

@@ 2.3.2 Divide-and-conquer algorithm
With a metric to quantify the CP structure of a network, we can present the divide-and-conquer algorithm. We slightly modify the algorithm in Yanchenko, (2022) for reasons that will be explained later. The idea is to find the CP structure on small sub-networks of the network and then combine these results together to yield the CP labels on the entire network. Specifically, let $G$ be the graph with $n$ nodes, $q\in(0,1)$ be the proportion of nodes to sub-sample and $B$ the number of sub-sample to draw. First, we sample a sub-graph with $qn$ nodes using some graph sub-sampling routine, and find the optimal CP labels from (3). We repeat this process $B$ times and output $\hat{\boldsymbol{c}}\in[0,1]^{n}$ where $\hat{c}_{i}$ is the number of times node $i$ was assigned to the core divided by the total number of samples $B$. The full algorithm is presented in Algorithm 2.
@@
Input: graph $G$, proportion of nodes to sub-sample $q\in(0,1)$, number of sub-samples $B\in\mathbb{Z}^{+}$, sub-sampling scheme $\mathcal{S}$
@@
       Randomly sample $G$ using $\mathcal{S}$ to obtain sub-network $S^{(b)}$  Obtain CP labels of $S^{(b)}$, $\hat{\boldsymbol{c}}^{(b)}\in\{0,1\}^{qn}$
@@
  Obtain $\hat{\boldsymbol{c}}$ as $\hat{c}_{i}=\sum_{b=1}^{B}\hat{c}_{i}^{(b)}/B$
@@
This algorithm is slightly different from the original paper. Originally, the output was the number of times node $i$ was assigned to the core divided by the number of times that node $i$ was sampled. Indeed, this approach more closely resembles the stitching routine of the $\mathsf{PACE}$ algorithm. The reason that $\mathsf{PACE}$ divides by the number of times that nodes $i$ and $j$ were sub-sampled together, $N_{ij}$, and then only keeps the estimate if $N_{ij}>\beta$ is that the estimate $\hat{C}_{ij}$ would be unreliable if it was constructed from only a few sub-samples. For the CP task, however, if a node is sub-sampled infrequently, this gives us more information, rather than less. Indeed, core nodes are those that are more central or influential in the network. Therefore, if a node is sampled infrequently, then it is unlikely to be important and also unlikely to be a part of the core. Thus, dividing by the total number of samples will decrease the estimate of $\hat{c}_{i}$ for infrequently sampled nodes. Implicitly, this means we should favor sub-sampling methods which have a high probability of sampling core nodes. As we will see, this small change not only makes proving theoretical statements easier, but also improves the empirical performance of the algorithm.
@@ 2.4 Sub-graph sampling
The simplest method is to randomly sample nodes without replacement. Each node has an equal probability of being selected, and once all $qn$ nodes are chosen, then the edges between these nodes make up the sub-graph. This method was originally considered in both Mukherjee et al., (2021) and Yanchenko, (2022). While this is the simplest sampling scheme, it is unclear if it will preserve important properties of the network. For example, Richardson et al., (2003) show that if the degrees of the full network have a power-law distribution, then RN sampling will not necessarily yield the same degree distribution on the sub-graphs.
@@
In this method, nodes are again sampled at random, but now their probability of being sampled is proportional to their degree. Once the $qn$ nodes are sampled, then the sub-graph is completed with all associated edges between sampled nodes. Mathematically, this means sampling without replacement from a distribution with probability mass function $\mathsf{P}(Z=j)=d_{j}/\sum_{k=1}^{n}d_{k}$ where $d_{j}$ is the degree of node $j\in\{1,2,\dots,n\}$. As Leskovec and Faloutsos, (2006) point out, this approach will likely lead to sub-graphs which have a different degree distribution from the original graph, while also being denser.
@@
Instead of randomly sampling nodes, this approach samples edges uniformly at random. Edges are drawn with equal probability and the two nodes connected by the sampled edge are added to the sub-graph. This process continues until $qn$ nodes have been sampled, at which point the remaining edges between sampled nodes are also added. Adding the remaining edges between sampled nodes is sometimes called the Induction Step (Ahmed et al.,, 2013) and will be used in all remaining algorithms as well. While edges can be selected with equal probability, if the graph has edge weights, then these could be used for a sampling scheme analogous to DN.
@@
The previous three methods randomly sampled nodes or edges. For the remaining methods, we will consider graph exploration approaches. The BFS algorithm begins by sampling a node at random. Then all neighbors of this sampled node are included in the sub-graph. Next, all neighbors of the neighbors are also included in the sub-sample. This process continues until we have sampled $qn$ nodes.
@@
DFS can be considered as the complement of a BFS algorithm. This method also begins by randomly sampling a starting node. Then instead of including all neighbors of this starting node, it randomly chooses one neighbor. From there, it visits a neighboring node which has yet to be visited. This process continues until a node is reached such that all of its neighbors have already been visited. At this point, the algorithm goes back to the nearest unvisited node and continues the process again. The process continues until $qn$ nodes have been sampled. Krishnamurthy et al., (2005) studied the ability of BFS and DFS to preserve certain graph properties in the sub-samples, but found both performed poorly.
@@
This approach shares similarities to BFS. As in BFS, a node is sampled at random and then all of its neighbors are also added to the sub-graph. Instead of looking at the neighbors of neighbors as in BFS, RNN samples a new node and includes all of the new node’s neighbors. This continues until $qn$ nodes have been sampled. This approach is equivalent to sampling $1$-hop neighborhoods (e.g. Mukherjee et al.,, 2021), and similar to an ego neighborhood except that the root node is also included in the sub-sample.
@@
In our final algorithm, we traverse the network through a random walk. Specifically, we begin by sampling a node uniformly at random. Then from this node, one of its neighbors is sampled at random, and this process continues until the walk has traversed $qn$ nodes. If the walk ends before reaching $qn$ unique nodes, then we restart the walk from a randomly selected node. Various works have studied the properties of RW-based sub-graphs (Ribeiro et al.,, 2012; Li et al.,, 2015; Chiericetti et al.,, 2016). In particular, Lovász, (1993) shows that RW yields a uniform distribution on the edges, implying that RE and RW should perform similarly.
@@ 3 Theoretical Results
In this section, we present theoretical results for each algorithm. As this paper’s focus is on the sub-sampling step, we will primarily focus on the role of the sampling scheme on the theoretical results. Additionally, Mukherjee et al., (2021) devoted significant attention to the theoretical properties of $\mathsf{PACE}$, so we only briefly discuss these before developing new theoretical results for the CP divide-and-conquer algorithm.
@@ 3.1 $\mathsf{PACE}$ algorithm
The goal of the theoretical results for $\mathsf{PACE}$ is to show that the estimated clustering matrix $\hat{\bf C}$, is “close” to the true clustering matrix ${\bf C}$ which generated the network. Recall that $C_{ij}=1$, if nodes $i$ and $j$ are in the same community, and 0 otherwise. To formalize this notion, let $\tilde{\delta}({\bf C},\hat{\bf C})$ be the difference between the true ${\bf C}$ and estimate $\hat{\bf C}$ where
$\tilde{\delta}({\bf C},{\bf C}^{\prime})=\frac{1}{n^{2}}||{\bf C}-{\bf C}^{\prime}||_{F}^{2}$
and $||\cdot||_{F}$ is the Frobenius norm. Thus, $\tilde{\delta}({\bf C},\hat{\bf C})$ is the mis-clustering rate of the $\mathsf{PACE}$ algorithm. In the following theorem, the authors bound the expectation of this term under a general community detection algorithm and sampling scheme.
Theorem 3.1 (Theorem 3.1 from Mukherjee et al., (2021)): Let graph $G$ have $K$ communities where community $k\in\{1,\dots,K\}$ contains $p_{k}n$ nodes, and define $p_{max}=\max\{p_{1},\dots,p_{K}\}$ as the proportion of nodes in the largest community. Given some sampling scheme $\mathcal{S}$, let $S$ be a randomly chosen sub-graph of the network. Moroever, let $\hat{\bf C}$ be the estimated clustering matrix returned by $\mathsf{PACE}$. Then the expected mis-clustering rate, $\mathbb{E}\tilde{\delta}({\bf C},\hat{\bf C})$ can be bounded by:
$\mathbb{E}\tilde{\delta}({\bf C},\hat{\bf C})\leq\frac{B}{\beta n^{2}}\mathbb{E}||\hat{\bf C}^{(S)}-{\bf C}^{(S)}||^{2}_{F}+p_{max}\max_{i,j}\mathbb{P}(N_{ij}<\beta)$ (4)
where $\hat{\bf C}^{(S)}$ and ${\bf C}^{(S)}$ are the estimated and true clustering matrices, respectively, restricted to sub-graph $S$.
There are two main parts to the bound in (4). The first term quantifies the performance of the community detection algorithm on a randomly chosen sub-graph $S$ which is mainly a function of the clustering algorithm. On the other hand, the second term measures how well the graph is covered by the sub-samples and depends only on the sub-sampling procedure.
As our interest is on the sub-sampling step, we focus on this second term. Indeed, the authors of $\mathsf{PACE}$ provide two special cases of the results under the RN sampling scheme and ego neighborhood sampling. As we do not discuss ego neighborhoods, we omit discussion of this result. The following gives the results for RN sampling.
Corollary 3.1 (Corollary 3.1 from Mukherjee et al., (2021)): Assume that RN is used to construct the subgraphs. Then
@@
This result shows that the mis-clustering error coming from the sub-sampling step decreases exponentially with $B$, the total number of sub-samples drawn, as well as $q$, the proportion of nodes in each sub-graph. As the authors note, deriving the specific result for other sub-sampling routines is difficult since the distribution of $N_{ij}$ is not simply $N_{ij}\sim\mathsf{Binomial}(B,q)$ as in the RN setting.
@@ 3.2 CP algorithm
Using similar ideas to that of Mukherjee et al., (2021), we present novel theoretical results for the CP divide-and-conquer algorithm. As in the previous sub-section, we are interested in bounding the expected mis-classification error.
@@ 3.2.1 General mis-classification result
First, we derive the expression for the general mis-classification rate. Recall from Algorithm 2, our estimate $\hat{\boldsymbol{c}}$ is defined as
@@
where $\hat{\boldsymbol{c}}^{(b)}$ are the CP labels from sub-sample $b$, i.e., $\hat{c}_{i}^{(b)}=1$ if node $i$ was assigned to the core in sub-sample $b$, and 0 otherwise. Then we define the mis-clustering of the algorithm as
@@
where $\boldsymbol{c}^{*}$ are the true CP labels, i.e., $c_{i}^{*}=1$ if node $i$ is a core node in the data-generating process, and 0 otherwise. The following result is analogous to Theorem 3.1 from Mukherjee et al., (2021).
Theorem 3.2: Given some graph $G$ with adjacency matrix ${\bf A}$ and sampling scheme $\mathcal{S}$, let $S$ be a randomly chosen sub-graph. Let $\hat{\boldsymbol{c}}$ be the estimated CP labels using the CP divide-and-conquer algorithm. Then the expected mis-clustering rate, $\mathbb{E}\delta(\hat{\boldsymbol{c}},\boldsymbol{c}^{*})$ can be bounded by
@@
Please see the Supplemental Materials for all proofs in this sub-section. Similar to the results of $\mathsf{PACE}$, our error bound has two terms; the first which primarily depends on the CP algorithm, and the second which only depends on the sub-sampling routine. Indeed, if the expected fraction of mis-classified nodes and the expected fraction of core nodes that are not sampled both go to 0 for large $n$, then this entire term will converge to 0. One noticeable difference between the two results is that Theorem 3.2 does not depend on $B$, the number of sub-samples, while Theorem 3.1 does. This discrepancy is because the CP divide-and-conquer algorithm divides the number of times a node was assigned to the core by $B$, rather than the number of times it was sub-sampled. If we had taken the latter approach, then $B$ would have appeared in the final result, along with a term counting the number of times a node was sampled.
@@ 3.2.2 Specific sub-sampling results
The results of Theorem 3.2 are agnostic to the network-generating model, CP detection algorithm and sub-sampling scheme. By assuming a particular model, we can find the analytic expression for the second term in (9) for the several of the sub-sampling schemes in Section 2.4.
@@
We assume networks are generated with a CP structure from the stochastic block model (SBM) (Holland et al.,, 1983). For adjacency matrix ${\bf A}$, let $A_{ij}\stackrel{{\scriptstyle ind.}}{{\sim}}\mathsf{Bernoulli}(\varrho_{n}P_{ij})$ where $P$ corresponds to a SBM that does not depend on $n$, and $\varrho_{n}=o(1)$ is a sparsity-inducing parameter. Furthermore, if $\boldsymbol{c}^{*}$ correspond to the true CP labels, then
$P_{ij}=\begin{cases}p_{11}&\text{if }c_{i}^{*}=c_{j}^{*}=1\\ p_{22}&\text{if }c_{i}^{*}=c_{j}^{*}=0\\ p_{12}&\text{otherwise}.\\ \end{cases}$
@@
If $p_{11}>p_{12}>p_{22}$, then this model generates networks with a CP structure, so we call it the CP-SBM. We may also assume that $\alpha_{n}=k/n$, the proportion of core nodes, is such that $\alpha_{n}=o(1)$ to model the empirical observation that the core is typically smaller than the periphery. In the following corollary, we find an analytic expression for the expected number of core nodes that are not sampled in each sub-graph, the second term in (9).
@@
Corollary 3.2: Let ${\bf A}$ be generated from a CP-SBM with $n$ nodes and $k$ core nodes. Let $S$ be a randomly chosen sub-graph of ${\bf A}$ of size $q_{n}n$ where $q_{n}\in(0,1)$ may depend on $n$. Furthermore, let $y_{i}^{(S)}$ is the event that node $i$ is a core node and not sampled in $S$. If RN is used to sample $S$ and $q_{n}=O(1)$, then
$\mathbb{E}\frac{1}{n}\sum_{i=1}^{n}y_{i}^{(S)}=(1-q)\frac{k}{n}.$ (7)
If $S$ is sampled via DN or RE and $q_{n}=o(1)$, then
$\lim_{n\to\infty}\mathbb{E}\frac{1}{n}\sum_{i=1}^{n}y_{i}^{(S)}=\left(1-q_{n}\frac{p_{12}}{p_{22}}\right)\alpha_{n}.$ (8)
The proof carefully computes the expected number of cores nodes that are sampled in each sub-graph. For RN, the number of core nodes sampled is exactly hyper-geometric distributed, meaning the result is non-asymptotic and does not require any assumptions on $q$. For DN and RE on the other hand, we are forced to rely on the node’s expected degree, leading to the asymptotic result. The assumption $q_{n}=o(1)$ is needed to ensure that the probability of sampling the same node twice is negligible. It also ensures that there exists some $N$ such that for all $n\geq N$, $q_{n}\tfrac{p_{12}}{p_{22}}<1$ such that the expectation is always positive. Assuming $q_{n}=o(1)$ is reasonable as otherwise the number of nodes in the sub-sample would grow linearly with $n$. Given this assumption, the sub-sample size can still grow with $n$ but it must be sub-linear.
@@
In both (11) and (12), as size of the sub-graphs increases ($q_{n}$), this term decreases. Moreover, this expression depends on $\alpha_{n}=k/n$. Assuming that the periphery dominates the core, this entire term will go to zero regardless of the sampling scheme which theoretically validates our decision to divide the core estimates by $B$. Of course, the sub-sampling routine will still affect the performance for finite $n$.
@@
It is also noteworthy that DN and RE yield identical expressions. This means these sampling schemes are equivalent in terms of the number of core nodes that are expected to be sampled in each sub-graph when the network is generated from a CP-SBM. This does not ensure that they will perform identically as the sub-sampling scheme affects the topology of the sub-graph $S$ which may affect the first term in (9). Moreover, in a CP-SBM we have $p_{12}>p_{22}\implies p_{12}/p_{22}>1$, meaning that DN and RE have a uniformly smaller expected error than RN. Additionally, as the strength of the CP structure of the network increases, the error will decrease because $p_{12}/p_{22}$ increases. In other words, we expect a lower error rate if the network exhibits a stronger CP structure.
@@
It is important to highlight that these results and discussion apply only to the second term in (9). It is possible that a sub-sampling method may not sample core nodes as often, but has a lower mis-classification rate on the sub-graph, leading to a smaller value of the first term in (9), or vice-versa. Finally, we briefly mention why we have not derived results for the other sub-sampler algorithms. For RNN, the difficulty is that the number of nodes sampled on each draw is different since we are sampling a node and its neighbors. BFS and DFS depend more heavily on the observed graph as opposed to the expected behavior of the nodes, complicating the analysis. Lastly, it is difficult to keep track of the visited nodes in RW. Deriving the results for these sub-sampling routines is another important avenue of future work.